# Algebraic K-theory of finite algebras over higher local fields

### v2

@@ 4. Segal conjecture
[…]
@@
Let $F$ be any $p$-local type $n+2$ finite complex, if $\mathbb{F}_{p}\otimes W\not\simeq 0$, where $W$ is the fiber of $\varphi:F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)\to F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}$. Then $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[\Omega S^{3}])$ is not bounded, and also the $F$-Segal conjecture fails for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[\Omega S^{3}])$. In fact, $\mathbb{F}_{p}\otimes W\not\simeq 0$.
@@
By [BnrM22, §5, Lemma 4.9, Lemma 5.5], $\operatorname{THH}(\mathbb{S}[\Omega S^{3}])\simeq\bigoplus_{m\geq 0}B_{m}$, and then condition (1), (2), (3) in Proposition 4.27 hold.  [^6] Actually, note that as $\mathbb{E}_{1}$-algebras $\mathbb{S}[\Omega S^{3}]$ is equivalent to $\oplus_{k\geq 0}(S^{2})^{\otimes k}$. By [Mat21, Theorem 3.8], we know that $\operatorname{THH}(\mathbb{S}[\Omega S^{3}])\simeq\bigoplus_{m\geq 0}\textup{Ind}^{S^{1}}_{C_{m}}(S^{2})^{\otimes m}$ in $\operatorname{Fun}(\textup{B}S^{1},\operatorname{Sp})$. So the decompositions seems equivalent. First, we claim following fact.
@@
This map induce equivalence on homotopy groups, by [BnrM22, Lemma 5.5], and the fact that $\textup{Ind}^{S^{1}}_{C_{m}}S^{2m}\simeq S^{2m}\oplus\Sigma S^{2m}$. Now both sides are connective spectra, then we conclude $B_{m}\simeq\textup{Ind}^{S^{1}}_{C_{m}}S^{2m}\simeq S^{2m}\oplus S^{2m+1}$ (this can be seen by filtering $\mathbb{S}$ as $\tau_{\geq*}\mathbb{S}$, and consider the graded). ∎
@@
Now, we need to show that $\mathbb{S}/p\otimes W$ is not 00. Suppose $\mathbb{S}/p\otimes W\simeq 0$, then $\mathbb{F}_{p}\otimes W\simeq\mathbb{Z}\otimes\mathbb{S}/p\otimes W\simeq 0$. But this will contradicts to following fact, therefore $\mathbb{S}/p\otimes W\not\simeq 0$.
@@
However, when we replace $\operatorname{BP}\langle n\rangle$ with $\operatorname{BP}$ the Segal conjecture holds, and even for $\operatorname{MU}$. First recall that in [LeNR11, Theorem 1.1], the authors showed that the following Frobenius maps are $p$-adic equivalences.
@@
where the second map is an equivalence. The fiber of $\mathbb{S}/p\otimes\operatorname{THH}(\operatorname{BP})\to\mathbb{S}/p\otimes\operatorname{THH}(\operatorname{BP})^{tC_{p}}$ is 00. Hence the first map induce isomorphism on homotopy groups. Thus the total Frobenius map induces isomorphism on homotopy groups. ∎
@@
$\operatorname{THH}(\operatorname{BP}[x]/x^{e})\to\operatorname{THH}(\operatorname{BP}[x]/x^{e}),$
$\operatorname{THH}(\operatorname{MU}[x]/x^{e})\to\operatorname{THH}(\operatorname{MU}[x]/x^{e}).$
@@
We only prove the cases for $\operatorname{BP}$, according to Lemma 4.21, the Frobenius map is a product of the following map
@@
where the second map is an equivalence. After tensoring with $\mathbb{S}/p$ the fist map has fiber equivalent to 00 by [LeNR11, Theorem 1.1], then the Frobenius map is a $p$-adic equivalence. [^7] Note that, using the above argument, if one replaces $\operatorname{THH}(\operatorname{BP})$ with $\operatorname{THH}(\mathbb{S})\simeq\mathbb{S}$, then one proves that $\operatorname{THH}(\mathbb{S}[x]/x^{e})\to\operatorname{THH}(\mathbb{S}[x]/x^{e})^{tC_{p}}$ is an $p$-adic equivalence.
@@
$\operatorname{THH}_{k}(\mathbb{F}_{p}\otimes\mathbb{S}[v_{0}]\otimes\cdots\otimes\mathbb{S}[v_{n-1}]\otimes\mathbb{S}[v_{n}]/v_{n}^{e}])\to\operatorname{THH}_{k}(\mathbb{F}_{p}\otimes\mathbb{S}[v_{0}]\otimes\cdots\otimes\mathbb{S}[v_{n-1}]\otimes\mathbb{S}[v_{n}]/v_{n}^{e}])^{tC_{p}}$
@@
Because $\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ is connective, then the summand $\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes\operatorname{Hom}((S^{1}/C_{k})_{+},(\Sigma^{2}\operatorname{BP}\langle n\rangle)^{\otimes k})$ is $12k-$ 1 connected, so when $k\to\infty$, the connectivity of the summand tends to infinity. Thus we have the following equivalences.
@@ 5. Descent spectral sequence
Note that $\operatorname{THH}(\operatorname{MU}[x])\to\operatorname{MU}[x]$ is evenly free. We need to show that for any map of $\mathbb{E}_{1}$-rings $\operatorname{THH}(\operatorname{MU}[x])\to A$, where $A$ is even, the push-out $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is even and free over $A$. Recall that $\operatorname{MU}[x]\otimes\operatorname{MU}[x]$ has homotopy as $(\operatorname{MU}_{*}\operatorname{MU})[x,y]\simeq\operatorname{MU}_{*}[x,x^{\prime}][x_{i}|i\in\mathbb{N}]$. Therefore $\operatorname{THH}(\operatorname{MU}[x])_{*}\simeq\Lambda_{\operatorname{MU}_{*}[x]}(\sigma x_{i},\sigma x)$, which is odd. Hence the classes $\sigma x_{i},\sigma x$ maps to 00 in $\pi_{*}A$, therefore the Tor spectral sequence computes $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is $A_{*}\otimes_{\operatorname{MU}_{*}[x]}\Gamma_{\operatorname{MU}_{*}[x]}(\sigma^{2}x_{i},\sigma^{2}x;i\in\mathbb{N})$ which is concentrated on even degrees, hence the push-out $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is even. Now, combining with Proposition 2.32, we have following fact.
@@
In general, if $A$ is an $\mathbb{E}_{3}$-algebra, and $B$ is an $\mathbb{E}_{2}$-$A$-algebra, then we could view $A,B$ as filtered algebras via Whitehead tower. And form the cosimplicial filtered object
$[n]\mapsto B^{\otimes_{A}[n]},$
@@ 6. Computation for height one case
In height $-1$ case, the computation of $\operatorname{TC}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ was presented in [HM97a]. We will use the strategy and the computations in [AR02] to carry out some computations for $\operatorname{TC}(\ell[x]/x^{e})$, where we use $\ell$ to denote $\operatorname{BP}\langle 1\rangle$.
@@
$\bigoplus_{k=3}^{2n+2}\Lambda(u_{n+1},\lambda_{[k]}^{\prime})\otimes P_{r(k-2)}(t\mu)\otimes\mathbb{F}_{p}\{\lambda_{[k]}t^{i}|v_{p}(i)=k-1\}.$
The associated graded of the homotopy fixed pointed spectral sequence computing $V(1)_{*}(\operatorname{THH}(\ell)^{hC_{p^{n}}})$ maps by a $(2p-2)$-coconnected map to
@@
We know that $v_{2}$ is detected by $t\mu$ in $V(1)_{*}\operatorname{TC}^{-}(\ell)$. Let $A=\Lambda_{\mathbb{F}_{p}}(\lambda_{1},\lambda_{2})\otimes P(t^{p^{2n+2}})$, $B=\bigoplus_{k=3}^{2n+2}\Lambda(u_{n+1},\lambda_{[k]}^{\prime})\otimes\mathbb{F}_{p}\{\lambda_{[k]}t^{i}|v_{p}(i)=k-1\}$, then $\hat{E}^{\infty}(C_{p^{n+1}})/(t\mu)=A\oplus B$. We know that both canonical map and Frobenius map send $\lambda_{1},\lambda_{2}$ to the same name elements in the target. Hence these classes will contribute to the homotopy groups on $V(1)_{*}\operatorname{TC}(\ell[x]/x^{e})/(v_{2})$, i.e. we have following fact.

### v3

@@ 4. Segal conjecture
${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle[\Omega Y])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}\scriptstyle{\simeq}$${F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle[\Omega Y])^{tC_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\prod_{m\geq 0}F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m/p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{\prime}}$$\scriptstyle{\operatorname{id}\otimes\varphi\otimes\varphi_{m}}$${F\otimes\prod_{m\geq 0}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\prod_{m\geq 0}F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\otimes B_{m}^{tC_{p}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${\prod_{m\geq 0}F\otimes(\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes B_{m})^{tC_{p}}.}$
@@
Let $F$ be any $p$-local type $n+2$ finite complex. Let $W$ be the fiber of $\varphi:F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)\to F\otimes\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}$. Then the following statements are true.
  1. (1)
    If $\mathbb{S}/p\otimes W\not\simeq 0$, then $F\otimes\operatorname{TR}(\operatorname{BP}\langle n\rangle[\Omega S^{3}])$ is not bounded, and also the $F$-Segal conjecture fails for $\operatorname{THH}(\operatorname{BP}\langle n\rangle[\Omega S^{3}])$.
  2. (2)
    We have $\mathbb{F}_{p}\otimes W\neq 0$. In particular, $\mathbb{S}/p\otimes W\neq 0$.
@@
By [BM22, §5, Lemma 4.9, Lemma 5.5], $\operatorname{THH}(\mathbb{S}[\Omega S^{3}])\simeq\bigoplus_{m\geq 0}B_{m}$, and then condition (1), (2), (3) in Proposition 4.27 hold.  [^7] Actually, note that as $\mathbb{E}_{1}$-algebras $\mathbb{S}[\Omega S^{3}]$ is equivalent to $\oplus_{k\geq 0}(S^{2})^{\otimes k}$. By [MAT21, Theorem 3.8], we know that $\operatorname{THH}(\mathbb{S}[\Omega S^{3}])\simeq\bigoplus_{m\geq 0}\textup{Ind}^{S^{1}}_{C_{m}}(S^{2})^{\otimes m}$ in $\operatorname{Fun}(\textup{B}S^{1},\operatorname{Sp})$. So the decompositions seems equivalent. First, we claim following fact.
@@
This map induce equivalence on homotopy groups, by [BM22, Lemma 5.5], and the fact that $\textup{Ind}^{S^{1}}_{C_{m}}S^{2m}\simeq S^{2m}\oplus\Sigma S^{2m}$. Now both sides are connective spectra, then we conclude $B_{m}\simeq\textup{Ind}^{S^{1}}_{C_{m}}S^{2m}\simeq S^{2m}\oplus S^{2m+1}$ (this can be seen by filtering $\mathbb{S}$ as $\tau_{\geq*}\mathbb{S}$, and consider the graded). ∎
@@
Now, we need to show that $\mathbb{S}/p\otimes W$ is not 0. Suppose $\mathbb{S}/p\otimes W\simeq 0$, then $\mathbb{F}_{p}\otimes W\simeq\mathbb{Z}\otimes\mathbb{S}/p\otimes W\simeq 0$. But this will contradicts to following fact, therefore $\mathbb{S}/p\otimes W\not\simeq 0$.
@@
However, when we replace $\operatorname{BP}\langle n\rangle$ with $\operatorname{BP}$ the Segal conjecture holds, and even for $\operatorname{MU}$. First recall that in [LR11, Theorem 1.1], the authors showed that the following Frobenius maps are $p$-adic equivalences.
@@
where the second map is an equivalence. The fiber of $\mathbb{S}/p\otimes\operatorname{THH}(\operatorname{BP})\to\mathbb{S}/p\otimes\operatorname{THH}(\operatorname{BP})^{tC_{p}}$ is 0. Hence the first map induce isomorphism on homotopy groups. Thus the total Frobenius map induces isomorphism on homotopy groups. ∎
@@
$\operatorname{THH}(\operatorname{BP}[x]/x^{e})\to\operatorname{THH}(\operatorname{BP}[x]/x^{e})^{tC_{p}},$
$\operatorname{THH}(\operatorname{MU}[x]/x^{e})\to\operatorname{THH}(\operatorname{MU}[x]/x^{e})^{tC_{p}}.$
@@
We only prove the cases for $\operatorname{BP}$. We can prove the Frobenius map is a product of the following map
@@
using a similar argument in Lemma 4. Here, the second map is an equivalence. After tensoring with $\mathbb{S}/p$ the fist map has fiber equivalent to 0 by [LR11, Theorem 1.1], then the Frobenius map is a $p$-adic equivalence. [^8] Note that, using the above argument, if one replaces $\operatorname{THH}(\operatorname{BP})$ with $\operatorname{THH}(\mathbb{S})\simeq\mathbb{S}$, then one proves that $\operatorname{THH}(\mathbb{S}[x]/x^{e})\to\operatorname{THH}(\mathbb{S}[x]/x^{e})^{tC_{p}}$ is an $p$-adic equivalence.
@@
$\operatorname{THH}_{k}(\mathbb{F}_{p}\otimes\mathbb{S}[v_{0}]\otimes\cdots\otimes\mathbb{S}[v_{n-1}]\otimes\mathbb{S}[v_{n}]/v_{n}^{e})\to\operatorname{THH}_{k}(\mathbb{F}_{p}\otimes\mathbb{S}[v_{0}]\otimes\cdots\otimes\mathbb{S}[v_{n-1}]\otimes\mathbb{S}[v_{n}]/v_{n}^{e})^{tC_{p}}$
@@
Because $\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ is connective, then the summand $\operatorname{THH}(\operatorname{BP}\langle n\rangle)\otimes\operatorname{Hom}((S^{1}/C_{k})_{+},(\Sigma^{2}\operatorname{BP}\langle n\rangle)^{\otimes k})$ is 2k12k-1 connected, so when $k\to\infty$, the connectivity of the summand tends to infinity. Thus we have the following equivalences.
@@ 5. Descent spectral sequence
Note that $\operatorname{THH}(\operatorname{MU}[x])\to\operatorname{MU}[x]$ is evenly free. We need to show that for any map of $\mathbb{E}_{1}$-rings $\operatorname{THH}(\operatorname{MU}[x])\to A$, where $A$ is even, the push-out $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is even and free over $A$. Recall that $\operatorname{MU}[x]\otimes\operatorname{MU}[x]$ has homotopy as $(\operatorname{MU}_{*}\operatorname{MU})[x,y]\simeq\operatorname{MU}_{*}[x,x^{\prime}][x_{i}|i\in\mathbb{N}]$. Therefore $\operatorname{THH}(\operatorname{MU}[x])_{*}\simeq\Lambda_{\operatorname{MU}_{*}[x]}(\sigma x_{i},\sigma x)$, which is odd. Hence the classes $\sigma x_{i},\sigma x$ maps to 0 in $\pi_{*}A$, therefore the Tor spectral sequence computes $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is $A_{*}\otimes_{\operatorname{MU}_{*}[x]}\Gamma_{\operatorname{MU}_{*}[x]}(\sigma^{2}x_{i},\sigma^{2}x;i\in\mathbb{N})$ which is concentrated on even degrees, hence the push-out $\operatorname{MU}[x]\otimes_{\operatorname{THH}(\operatorname{MU}[x])}A$ is even. Now, combining with Proposition 2.32, we have following fact.
@@
In general, if $A$ is an $\mathbb{E}_{3}$-algebra, and $B$ is an $\mathbb{E}_{2}$-AA-algebra, then we could view $A,B$ as filtered algebras via Whitehead tower. And form the cosimplicial filtered object
$n]\mapsto B^{\otimes_{A}[n]},$
@@ 6. Computation for height one case
In height 1-1 case, the computation of $\operatorname{TC}(\operatorname{BP}\langle n\rangle[x]/x^{e})$ was presented in [HM97a]. We will use the strategy and the computations in [AR02] to carry out some computations for $\operatorname{TC}(\ell[x]/x^{e})$, where we use $\ell$ to denote $\operatorname{BP}\langle 1\rangle$.
@@
$\bigoplus_{k=3}^{2n+2}\Lambda(u_{n+1},\lambda_{[k]}^{\prime})\otimes P_{r(k-2)}(t\mu)\otimes\mathbb{F}_{p}\{\lambda_{[k]}t^{i}\mid v_{p}(i)=k-1\}.$
The associated graded of the homotopy fixed pointed spectral sequence computing $V(1)_{*}(\operatorname{THH}(\ell)^{hC_{p^{n}}})$ maps by a (2p2)(2p-2)-coconnected map to
@@
We know that $v_{2}$ is detected by $t\mu$ in $V(1)_{*}\operatorname{TC}^{-}(\ell)$. Let $A=\Lambda_{\mathbb{F}_{p}}(\lambda_{1},\lambda_{2})\otimes P(t^{p^{2n+2}})$, $B=\bigoplus_{k=3}^{2n+2}\Lambda(u_{n+1},\lambda_{[k]}^{\prime})\otimes\mathbb{F}_{p}\{\lambda_{[k]}t^{i}\mid v_{p}(i)=k-1\}$, then $\hat{E}^{\infty}(C_{p^{n+1}})/(t\mu)=A\oplus B$. We know that both canonical map and Frobenius map send $\lambda_{1},\lambda_{2}$ to the same name elements in the target. Hence these classes will contribute to the homotopy groups on $V(1)_{*}\operatorname{TC}(\ell[x]/x^{e})/(v_{2})$, i.e. we have following fact.